A 1600kg van moving at 20m/s crashes into the back of a small car of mass 700kg traveling in the same direction at 15m/s.
(a) what is the velocity of the two vehicles immediately after the collision of the lock together?
(b) how much kinetic energy is transferred to other forms in the collision?
1
Expert's answer
2020-03-13T10:53:47-0400

a) From the conservation of momentum:


"Mv_1+mv_2=(m+M)u"

"1600(20)+700(15)=(700+1600)u"

"u=18.5\frac{m}{s}"

b)


"\Delta E=0.5Mv_1^2+0.5mv_2^2-0.5(m+M)u^2"

"\Delta E=0.5(1600)(20)^2+0.5(700)(15)^2-\\-0.5(700+1600)(18.5)^2=5160\ J"